# Interpretable Retinal Disease Prediction Using Biology-Informed Heterogeneous Graph Representations

### v1

@@ Abstract
Interpretability is crucial to enhance trust in machine learning models for medical diagnostics. However, most state-of-the-art image classifiers based on neural networks are not interpretable. As a result, clinicians often resort to known biomarkers for diagnosis, although biomarker-based classification typically performs worse than large neural networks. This work proposes a method that surpasses the performance of established machine learning models while simultaneously improving prediction interpretability for diabetic retinopathy staging from optical coherence tomography angiography (OCTA) images. Our method is based on a novel biology-informed heterogeneous graph representation that models retinal vessel segments, intercapillary areas, and the foveal avascular zone (FAZ) in a human-interpretable way. This graph representation allows us to frame diabetic retinopathy staging as a graph-level classification task, which we solve using an efficient graph neural network. We benchmark our method against well-established baselines, including classical biomarker-based classifiers, convolutional neural networks (CNNs), and vision transformers. Our model outperforms all baselines on two datasets. Crucially, we use our biology-informed graph to provide explanations of unprecedented detail. Our approach surpasses existing methods in precisely localizing and identifying critical vessels or intercapillary areas. In addition, we give informative and human-interpretable attributions to critical characteristics. Our work contributes to the development of clinical decision-support tools in ophthalmology.
I Introduction
Diabetic Retinopathy (DR), a complication of diabetes that affects the retinal vasculature, is one of the leading causes of blindness in adulthood [1]. It is associated with pathological changes to the retinal microvasculature, resulting in a widening of the intercapillary areas, and enlargement of the foveal avascular zone (FAZ). Currently, clinicians study biomarkers that capture these changes, such as blood vessel density (BVD), Fractal Dimension (FD), and FAZ area. These biomarkers can be extracted from retinal images, e.g., via optical coherence tomography angiography (OCTA), a non-invasive and high-resolution retinal imaging technique [2, 3]. OCTA enables the detailed examination of the retinal vasculature [4] down to the capillary level by resolving the superficial (SVC) and deep vascular complex (DVC), which enables early-stage DR detection [5].
@@ I Introduction
Current Machine Learning (ML)-based approaches trained on OCTA images and OCTA-extracted biomarkers have shown excellent performance for diagnosing DR [3, 6, 7, 8, 9, 10, 11]. However, these deep learning (DL) approaches are not used for clinical decision-making, partially because their outcomes are not interpretable, making it difficult for clinicians to verify or falsify model predictions. Although DL models may implicitly leverage high-level concepts (e.g., BVD or FAZ area), existing explainability methods fail to attribute the model’s decision to these. For convolutional neural networks (CNNs), CAM [12] and grad-CAM [13] are common methods that focus on localizing critical image regions. Previous studies [9, 10, 11] employed these explainability methods for DR classification with mixed results. While the approach allowed a rough identification of critical image regions, the emphasized areas remain broad and unspecific. Moreover, the most critical shortcoming of these methods is the semantic gap between the highlighted image regions and human-interpretable concepts. Unlike DL approaches, inherently interpretable models, such as decision trees based on human-understandable biomarkers, do not achieve state-of-the-art prediction accuracy.
@@
    We employ a graph learning architecture to our new representation for DR staging and exceed current state-of-the-art deep learning approaches and conventional biomarker-based solutions in prediction performance. With this application, we highlight the potential of our graph-based representation for downstream tasks.
@@
    Leveraging our heterogeneous graph representation, we introduce an explainability framework for DR staging that enables the precise identification of important structures in OCTA images and couples it with interpretable features at those specific locations.
@@ II Related Work
Figure 2: Processing pipeline for generating a heterogeneous graph that models an OCTA image’s most relevant biological concepts: vessels, intercapillary areas, and the FAZ. A vasculature segmentation [14] is used to create a vessel graph and intercapillary area graph and identify the FAZ. Finally, these components are merged into a single interconnected heterogeneous graph representation.
II-A Machine Learning for DR Staging in OCTA
@@
In recent years, the improvement of instrumentation [15] has made OCTA available to an increasing number of ophthalmologists. Since then, many studies have investigated DR with OCTA. Approaches to staging DR with OCTA can be broadly categorized into two groups. The first group comprises feature-based methods, which are based on prior knowledge of indicative biomarkers, the quantitative assessment of these markers, and decision-making based on the extracted information [3, 6, 7, 8]. Traditionally utilized biomarkers include BVD, blood vessel caliber (BVC), blood vessel tortuosity, vessel perimeter index, vessel complexity index, FAZ area, and FAZ contour irregularity. These biomarkers serve as inputs for classification models such as neural networks, support vector machines (SVM), or tree-based methods such as random forests (RF). Despite their widespread acceptance and proven significance for DR disease staging, feature-based approaches have limitations. They can only rely on a priori defined biomarkers, established based on human expertise and laborious verification procedures. In the classification setting, these biomarkers are usually evaluated as global metrics. Although localization of pathological changes is theoretically possible, it is not used for explanation generation in clinical decision support.
@@
The second group comprises DL models directly operating on the raw OCTA images. Previous studies adopted well-established CNN architectures such as ResNet [16], VGG [17], and EfficientNet [18] for the task and achieved high accuracy. Zhou et al. proposed a transformer-based foundation model for retinal images [19], which outperformed other state-of-the-art approaches for DR staging on color fundus images. Although the network was not directly trained on OCTA images, the model has a detailed understanding of abstract retinal concepts by training on a large amount of training data. DL-based methods have shown favorable performance compared to the aforementioned traditional approaches. This performance gap indicates that DL methods can extract information that is, to this point, not formalized as an established, known biomarker. They can be combined with explainability tools, such as CAM [12] and grad-CAM [13], to highlight critical image regions [9, 10, 11]. However, the human-in-the-loop cannot overcome the semantic gap between highlighted gray values and the abstract concepts established later in the CNN network. Clinicians cannot verify and interpret the algorithmic outcomes beyond the localization of influential regions. Moreover, the biomarkers utilized by the CNN networks cannot be identified and formalized.
II-B Graph Neural Networks as Explainable Models
@@
With the advent of graph neural networks (GNNs), convolutions were generalized towards graph-structured data; oftentimes including the concept of message-passing [20, 21, 22]. This has led to the widespread application of GNNs across diverse fields, including chemistry, physics, medicine, and social sciences, where graph-structured data is abundant [23]. In biomedical image analysis, successful applications of GNNs to discriminative tasks include brain disorder classification from functional magnetic resonance images [24], multiple sclerosis disease activity prediction from MR images [25], vessel hierarchy classification in microscopic images [26] and pathology image classification [27].
@@
In tandem with the evolution and utilization of GNN architectures, the methods for explaining their predictions advanced. For example, gradient-based approaches, proven successful in computer vision and CNNs, effectively provided explanations for graph-based molecular property prediction [28]. Specific perturbation-based techniques tailored for graph-structured data, such as GNNExplainer [29], have been developed. Notably, GraphLime [30], drawing inspiration from LIME [31], stands out as a surrogate-based explainability method that uses local perturbations. For a comprehensive overview of other gradient-, perturbation-, or surrogate-based explanation approaches, refer to [32]. Agarwal et al. [33] analyzed various explanation methods tailored for graph classification tasks. Their study showed that integrated gradients is the most faithful and accurate explanation method, particularly on real-world datasets with known ground truth explanations.
III Methods
III-A Construction of a Heterogeneous Graph Representation
@@ III-A1 Vessel Segmentation
We generate vessel segmentation maps using the method by Kreitner et al. [14, 34]. This method leverages synthetic training data to create high-fidelity label maps of the retinal vasculature. The resulting label maps include even the smallest capillary vessels and are highly continuous, which is important to generate continuous vessel and intercapillary area graphs in the next step.
III-A2 Vessel Graph
@@
Our representation incorporates vessel graphs [35, 36] to model the vasculature in imaged eyes. Biomarkers such as vessel complexity and diameter are also encapsulated in our vessel graphs. Each vessel segment between two bifurcation points is designated as a node, and edges between vessel nodes are established using bifurcation points. The vasculature segmentation serves as input for a graph extraction algorithm [37], which extracts the vessel graph and provides feature descriptors detailing properties such as volume, length, curvature, and surface-distance-based values for the centerline. These features serve as embeddings for the nodes in the vessel graph. For detailed descriptions of the extracted geometric features, refer to [37] and our github repository. Figure 3 depicts the vessel segments as nodes $v_{ves}$ and the edges between adjacent vessel segments as $e_{ves}$. We use the open-source Voreen framework, built to visualize and explore scientific data to run the graph extraction algorithm [38].
III-A3 Intercapillary Area Graph
@@
Intercapillary areas are indicators for disease progression in DR [39, 40]. Essentially, they are the counterparts of segmented vessels in OCTA images. We create the nodes of the intercapillary area graph using connected component labeling of the inverted segmentation map. The nodes of the intercapillary area graph are enriched with geometric descriptors such as total area, perimeter length, and eccentricity. We rely on the skeleton of the vessel segmentation to establish the edges in the intercapillary area graph. In the skeletonized segmentation map, we find edge candidates by traversing all skeleton pixels and checking if directly adjacent pixels are part of separate background components of the skeletonized segmentation map. Using the injective map of background components in the segmentation map to background components in the skeletonized segmentation map, edges are then introduced between the intercapillary areas. In Figure 3, the intercapillary area nodes $v_{\scriptscriptstyle ICA}$ represent the green areas (blue nodes), and neighborhood information is encoded in the $e_{\scriptscriptstyle ICA}$ edge type.
III-A4 Foveal avascular zone
@@
Because of its significance as a biomarker [41], we assign a separate node type and unique edge types to the FAZ. The FAZ is identified as the central intercapillary area of the connected component labeling. Our used datasets are FAZ-centered during preprocessing, note that this is a necessary prerequisite for correctly identifying the FAZ. The separate node type for the FAZ adds an inductive bias for downstream algorithms. The FAZ is represented through the node $v_{\scriptscriptstyle FAZ}$ (see Figure 3).
III-A5 Heterogeneous Edges
@@
Collectively, this biology-inspired graph representation introduces bias for downstream tasks while enhancing human interpretability by abstracting from gray values to intuitive concepts like vessels. Additionally, the graph representation preserves spatial information through positional coordinates and neighborhood information through edges. Positional coordinates can be embedded via x- and y-coordinates for intercapillary areas and FAZ, as well as x- and y-coordinates for the start and endpoints of vessel segments. Our graph representation can be used as a foundation for various prominent prediction tasks on OCTA images (see Section III-B and VI).
@@
III-B Graph Learning for DR Staging
We utilize our heterogeneous graph representation to predict the DR stage of the imaged eye, which we frame as a graph classification task. We use the Pytorch Geometric framework to implement our GNN model [42]. Specifically, we use a GNN architecture, as depicted in Fig. 4, tailored to our presented heterogeneous graph structure:
@@ III-B Graph Learning for DR Staging
    First, we preprocess the different node types separately using a block of two fully connected layers for each type. Each layer is followed by a batch norm layer for regularization and a ReLU activation function, see Fig 4 (1). These preprocessing layers allow to adjust the node embeddings for the following message-passing layer.
@@
    In the next step, we refine the embeddings based on neighborhood information using one independent SAGE [22] layer for each edge type, see Fig 4 (2). The update using one specific type of neighborhood node, e.g. type $a$ is then defined as
@@
    Notably, all our edges are bidirectional and allow message passing in both directions. This message-passing layer is used to refine the node embeddings based on the neighborhood information stored in the graph. Using heterogeneous edges, node embeddings can be refined using the information of all types of connected neighbors. The resulting embeddings from the refinement using distinct types of neighborhood nodes are then aggregated
@@
    After the graph layers, another block of two fully connected layers postprocesses the individual representations of the different node types, see Fig 4 (3). Again, batch norm layers and a ReLU activation function are applied following each layer.
@@
    Next, we use sum and max pooling to aggregate over the embeddings of all nodes of the same type. The results of both aggregations are then concatenated, followed by another concatenation of all aggregated embeddings of the different node types, see Fig 4 (4). Using a sparse aggregation (max) and a dense aggregation (sum) allows the network to extract global information and, at the same time, pay special attention to single instances.
@@
    Finally, we use an MLP head consisting of four fully connected layers with dropout on the aggregated fixed-size feature vector for multiclass classification, see Fig 4 (5).
Figure 4: Illustration of the graph neural network architecture designed for the DR disease staging task. The architecture takes a heterogeneous graph as input and predicts the disease stage for the provided input graph. We generalize homogeneous message-passing architectures to the heterogeneous case by learning individual message-passing functions for each node type. After the message passing layers and linear layers, an aggregation for each individual node type is performed. Finally, the embeddings for all node types are aggregated and passed on to a classification head.
Our proposed network is restricted in its complexity compared to current CNNs and transformer models. Table I shows a comparison of the number of parameters and the training time of the GNN model and other prominent network architectures. Notably, the graph construction step is a prerequisite for our approach. We observe that with non-runtime-optimized code the construction takes multiple seconds. Therefore, the benefit of faster training during hyperparameter tuning already outweighs the construction overhead.
TABLE I: Comparison of model complexities and training times for 100 epochs with a fixed batch size of 16 on a single NVIDIA RTX A6000. The heterogeneous GNN is more than 100 times smaller than the smallest evaluated CNN model. The * indicates that the model is fine-tuned without adapting all learnable weights. Pretraining required around two weeks with eight NVIDIA Tesla A100 [19].
@@
III-C Interpretability Framework for Retinal Graphs
Our explainability framework consists of two components. First, an attribution method, and second, a localization method. We combine these methods to enhance the interoperability of our graph-based classifier predictions.
III-C1 Attribution Method
We utilize integrated gradients [43, 44] as attribution method to describe the impact of specific nodes and their characteristics on the algorithmic outcome, following the evaluations by Agarwal et al. [33]. The integrated gradients approach follows a straight line path in the input feature space from a baseline sample to the target sample. In computer vision, this baseline sample is typically a zero-image. Equivalent to baseline images, graph-based approaches require a baseline vector for each node in the graph. The attribution score $a_{v_{i}}$ of a feature $i$ of a node $v$ given a baseline $x_{bl}$ is then defined as

### v2

@@ Abstract
Interpretability is crucial for utilizing machine learning models as clinical decision support tools for medical diagnostics. However, most state-of-the-art image classifiers based on neural networks are not interpretable. As a result, clinicians often resort to known biomarkers to guide diagnosis, although biomarker-based classification often suffers from drastic information loss compared to raw medical images. This work proposes a method that preserves the rich imaging information while simultaneously enhancing the interpretability of predictions for diabetic retinopathy staging from optical coherence tomography angiography (OCTA) images. Our method is based on a novel biology-informed heterogeneous graph representation that models retinal vessel segments, intercapillary areas, and the foveal avascular zone (FAZ) in a human-interpretable way. This graph representation allows us to frame diabetic retinopathy (DR) staging as a graph-level classification task, which we solve using an efficient graph neural network. We compare our method against established methods, including classical biomarker-based classifiers, convolutional neural networks (CNNs), and vision transformers in predicting the clinically assigned DR stage based on color fundus photography images. We find stage agreement rates of our method and alternative vision model based classifiers saturating at AUC-ROC values of 84%. Crucially, we use our biology-informed graph to provide explanations of unprecedented detail. Our approach surpasses existing methods in precisely localizing and identifying abnormal vessels and non-perfusion areas. Our approach sets the stage for the interpretable identification of patients who require special attention due to their traceable microvascular changes, only observable using the details of OCTA images.
[…]
organization=Munich Center for Machine Learning (MCML),city=Munich, country=Germany
[…]
organization=School of Medicine and Health, TUM University Hospital,city=Munich, country=Germany
[…]
organization=School of Computation, Information and Technology, TUM,city=Munich, country=Germany
[…]
organization=Weill Cornell Medicine, Cornell University, New York City, ,city=New York, country=USA
1 Introduction
Diabetic Retinopathy (DR), a complication of diabetes that affects the retinal vasculature, is one of the leading causes of blindness in adulthood [teo2021global]. It is associated with pathological changes to the retinal microvasculature, such as loss of capillaries through capillary degeneration, the occurrence of intraretinal microvascular abnormalities, microaneurysms, and also neovascularization.
Optical coherence tomography angiography (OCTA), a non-invasive and high-resolution retinal imaging technique [sun2019oct, le2021machine], has emerged as a powerful tool for studying these microvascular changes. OCTA enables detailed examination of the retinal vasculature [fayed2024retinal] down to the capillary level by resolving the superficial vascular complex (SVC) and deep vascular complex (DVC), facilitating early-stage DR detection [moore1999three].
Currently, clinicians study quantitative biomarkers extracted from OCTA images to characterize DR-related microvascular alterations, including vessel density, vessel tortuosity, and foveal avascular zone (FAZ) area. However, these biomarkers have important limitations: they fail to fully capture complex pathological changes such as intraretinal microvascular abnormalities and neovascularization, and simple metrics like tortuosity index and vessel density lack spatial localization, making clinical verification challenging. Deep learning models, that are already well-established for other imaging modalities [stolte2020survey], offer the potential to extract the full spectrum of microvascular pathological information from OCTA images to guide clinical diagnosis beyond traditional biomarkers. Such AI-based OCTA analysis could serve as a valuable support tool within the multimodal diagnostic workflow that integrates OCT, color fundus photography, fluorescein angiography, and patient history.
@@ 1 Introduction
Current Machine Learning (ML)-based approaches trained on OCTA images and OCTA-extracted biomarkers have shown good performance for diagnosing DR [le2021machine, alam2019supervised, sandhu2020automated, alam2020quantitative, heisler2020ensemble, ryu2021deep, zang2022diabetic]. However, these deep learning (DL) approaches are not used for clinical decision-making, partially because their black-box predictions cannot be effectively combined with the multi-modal information sources that ophthalmologists routinely use for DR assessment. Specifically, the lack of interpretability makes it difficult for clinicians to understand how OCTA-based predictions relate to findings from other imaging modalities and patient factors, hindering their integration into comprehensive diagnostic evaluations. Although DL models may implicitly leverage high-level pathological concepts (e.g., non-perfusion areas or microaneurysms), existing explainability methods fail to attribute the model’s decisions to these concepts. For convolutional neural networks (CNNs), CAM [zhou2016learning] and grad-CAM [selvaraju2017grad] are common methods that focus on localizing critical image regions. Previous studies [heisler2020ensemble, ryu2021deep, zang2022diabetic] employed these explainability methods for DR classification, yielding mixed results. While their approaches allowed a rough identification of critical image regions, the emphasized areas remain broad and unspecific. Moreover, the most critical shortcoming of these methods is the semantic gap between the highlighted image regions and human-interpretable concepts. Unlike DL approaches, inherently interpretable models, such as simple decision trees based on human-understandable biomarkers, do not achieve state-of-the-art prediction accuracy.
@@
    We employ a graph learning architecture to our new representation for DR staging that allows us to directly link stage predictions to clinically well-studied indicators of DR, such as abnormal vessel segments, microaneurysms, and retinal areas with decreased capillary density. With this application, we highlight the potential of our graph-based representation for further downstream tasks.
@@
    Finally, we utilize our interpretable model explanations combined with OCTA expert readings to identify a subgroup of patients where microvascular pathologies observed in 3 mm x 3 mm OCTA images deviate from the clinical DR stage based on color fundus images. This enables the identification of patients who require special attention for follow-up and treatment decisions.
@@ 2 Related Work
Figure 2: Processing pipeline for generating a heterogeneous graph that models an OCTA image’s most relevant biological concepts: vessels, intercapillary areas, and the FAZ. A vasculature segmentation [kreitner2023detailed] is used to create a vessel graph and intercapillary area graph and identify the FAZ. Finally, these components are merged into a single interconnected heterogeneous graph representation.
2.1 Machine Learning for DR Staging in OCTA
@@
In recent years, the improvement of instrumentation [kim2011vivo] has made OCTA available to an increasing number of ophthalmologists. Since then, many studies have investigated DR with OCTA. Approaches to staging DR with OCTA can be broadly categorized into two groups. The first group comprises feature-based methods, which are based on prior knowledge of indicative biomarkers, the quantitative assessment of these markers, and decision-making based on the extracted information [le2021machine, alam2019supervised, sandhu2020automated, alam2020quantitative]. Traditionally utilized biomarkers include blood vessel density (BVD), blood vessel caliber (BVC), blood vessel tortuosity, vessel perimeter index, vessel complexity index, FAZ area, and FAZ contour irregularity. These biomarkers serve as inputs for classification models such as neural networks, support vector machines (SVM), or tree-based methods such as random forests (RF). Despite their widespread acceptance and proven significance for DR disease staging, feature-based approaches have limitations. They can only rely on a priori defined biomarkers, established based on human expertise and laborious verification procedures. To date, no existing biomarker quantifies the severity of intraretinal microvascular abnormalities as well as the severity of microaneurysms in OCTA images. Moreover, in the classification setting, existing biomarkers are usually evaluated as global metrics without localization information.
@@
The second group comprises DL models directly operating on the raw OCTA images. Previous studies adopted well-established CNN architectures such as ResNet [xie2017aggregated], VGG [simonyan2014very], and EfficientNet [tan2019efficientnet] for DR staging and achieved high accuracy. Zhou et al. proposed a transformer-based foundation model for retinal images [zhou2023foundation], which outperformed other state-of-the-art approaches for DR staging on color fundus images. Although the network was not directly trained on OCTA images, the model has a detailed understanding of abstract retinal concepts by training on a large amount of data. DL-based methods have shown favorable performance compared to the aforementioned traditional approaches. This performance gap indicates that DL methods can extract information that, to date, has not been formalized as an established, known biomarker, such as the severity and frequency of intraretinal microvascular abnormalities. They can be combined with explainability tools, such as CAM [zhou2016learning] and grad-CAM [selvaraju2017grad], to highlight critical image regions [heisler2020ensemble, ryu2021deep, zang2022diabetic]. However, the human-in-the-loop cannot overcome the semantic gap between highlighted image regions and their corresponding intensity values and the abstract concepts established later in the CNN network. Clinicians cannot verify and interpret the algorithmic outcomes beyond the localization of influential regions. Moreover, the biomarkers utilized by the CNN networks cannot be identified and formalized.
2.2 Graph Neural Networks as Explainable Models
@@
With the advent of graph neural networks (GNNs), convolutions were generalized towards graph-structured data; oftentimes including the concept of message-passing [velickovic2017graph, kipf2016semi, hamilton2017inductive]. This has led to the widespread application of GNNs across diverse fields, including chemistry, physics, medicine, and social sciences, where graph-structured data is abundant [zhou2020graph]. In biomedical image analysis, successful applications of GNNs to discriminative tasks include brain disorder classification from functional magnetic resonance images [zhang2022classification], multiple sclerosis disease activity prediction from MR images [prabhakar2023self], vessel hierarchy classification in microscopic images [paetzold2021whole], and pathology image classification [wang2023ccf].
@@
In tandem with the evolution and utilization of GNN architectures, the methods for explaining their predictions have advanced. For example, gradient-based approaches, proven successful in computer vision and CNNs, effectively provided explanations for graph-based molecular property prediction [pope2019explainability]. Specific perturbation-based techniques tailored for graph-structured data, such as GNNExplainer [ying2019gnnexplainer], have been developed. Notably, GraphLime [huang2022graphlime], drawing inspiration from LIME [ribeiro2016should], stands out as a surrogate-based explainability method that uses local perturbations. For a comprehensive overview of other gradient-, perturbation-, or surrogate-based explanation approaches, refer to [yuan2022explainability]. Agarwal et al. [agarwal2023evaluating] analyzed various explanation methods tailored for graph classification tasks. Their study showed that integrated gradients is the most faithful and accurate explanation method, particularly on real-world datasets with known ground truth explanations.
3 Methods
3.1 Construction of a Heterogeneous Graph Representation
@@ 3.1.1 Vessel Segmentation
We generate vessel segmentation maps using the method by Kreitner et al. [kreitner2023detailed, menten2023synthetic]. This method leverages synthetic training data to create high-fidelity label maps of the retinal vasculature. The resulting label maps include even the smallest capillary vessels and are highly continuous, which is important to generate continuous vessel and intercapillary area graphs.
3.1.2 Vessel Graph
@@
Our representation incorporates vessel graphs [paetzold2023geometric, todorov2020machine] to model the vasculature in imaged eyes. Biomarkers such as vessel complexity and diameter are also encapsulated in our vessel graphs. Each vessel segment between two bifurcation points is designated as a node, and edges between vessel nodes are established using bifurcation points. The vasculature segmentation serves as input for a graph extraction algorithm [drees2021scalable], which extracts the vessel graph and provides feature descriptors detailing properties such as volume, length, curvature, and surface-distance-based values for the centerline. These features serve as embeddings for the nodes in the vessel graph. For detailed descriptions of the extracted geometric features, refer to [drees2021scalable] and our GitHub repository. Figure 3 depicts the vessel segments as nodes $v_{ves}$ and the edges between adjacent vessel segments as $e_{ves}$. We use the open-source Voreen framework, built to visualize and explore scientific data to run the graph extraction algorithm [meyer2009voreen].
3.1.3 Intercapillary Area Graph
@@
Intercapillary areas are indicators for disease progression in DR [schottenhamml2016automatic, terada2022intercapillary]. Essentially, they are the counterparts of segmented vessels in OCTA images. We create the nodes of the intercapillary area graph using connected component labeling of the inverted segmentation map. The nodes of the intercapillary area graph are enriched with geometric descriptors such as total area, perimeter length, and eccentricity. We rely on the skeleton of the vessel segmentation to establish the edges in the intercapillary area graph. In the skeletonized segmentation map, we find edge candidates by traversing all skeleton pixels and checking if directly adjacent pixels are part of separate background components of the skeletonized segmentation map. Using the injective map of background components in the segmentation map to background components in the skeletonized segmentation map, edges are then introduced between the intercapillary areas. In Figure 3, the intercapillary area nodes $v_{\scriptscriptstyle ICA}$ represent the green areas (blue nodes), and neighborhood information is encoded in the $e_{\scriptscriptstyle ICA}$ edge type.
3.1.4 Foveal avascular zone
@@
Because of its significance as a biomarker [sun2021optical, conrath2005foveal], we assign a separate node type and unique edge types to the FAZ. The FAZ is identified as the central intercapillary area of the connected component labeling. Our used datasets are FAZ-centered during preprocessing; note that this is a necessary prerequisite for correctly identifying the FAZ. The separate node type for the FAZ adds an inductive bias for downstream algorithms. The FAZ is represented through the node $v_{\scriptscriptstyle FAZ}$ (see Figure 3).
3.1.5 Heterogeneous Edges
@@
Collectively, this biology-inspired graph representation introduces bias for downstream tasks while enhancing human interpretability by abstracting from gray values to intuitive concepts like vessels. Additionally, the graph representation preserves spatial information through superpixels in the image and neighborhood information through edges. Positional coordinates can be embedded via x- and y-coordinates for intercapillary areas and FAZ, as well as x- and y-coordinates for the start and endpoints of vessel segments. Our graph representation can be used as a foundation for various prominent prediction tasks on OCTA images (see Section 3.2 and 6).
@@
3.2 Graph Learning for DR Staging
We utilize our heterogeneous graph representation to predict the DR stage of the imaged eye, which we frame as a graph classification task. We use the Pytorch Geometric framework to implement our GNN model [fey2019fast]. Specifically, we use a GNN architecture, as depicted in Fig. 4, tailored to our presented heterogeneous graph structure:
@@ 3.2 Graph Learning for DR Staging
    First, we preprocess the different node types separately using a block of fully connected layers for each type. Each layer is followed by a batch normalization layer for regularization and a ReLU activation function, as shown in Fig. 4 (1). These preprocessing layers can adjust the node embeddings for the following message-passing layer.
@@
    In the next step, we refine the embeddings based on neighborhood information using a block of independent message passing layers [hamilton2017inductive] for each edge type, see Fig 4 (2). The update using one specific type of neighborhood node, e.g., type $a$ is then defined as
@@
    All edges are bidirectional and allow message passing in both directions. This message-passing layer is used to refine the node embeddings based on the neighborhood information stored in the graph. Using heterogeneous edges, node embeddings can be refined using the information of all types of connected neighbors. The resulting embeddings from the refinement using distinct types of neighborhood nodes are then aggregated
@@
    After the graph layers, another block of fully connected layers postprocesses the individual representations of the different node types, see Fig 4 (3). Again, batch norm layers and a ReLU activation function are applied following each layer.
@@
    Next, global pooling is applied to aggregate over the embeddings of all nodes of the same type. All the aggregated embeddings of the different node types, see Fig 4 are then again aggregated into a single feature vector. (4).
@@
    Finally, we use an MLP head consisting of fully connected layers with dropout on the aggregated fixed-size feature vector for multiclass classification, see Fig 4 (5).
Figure 4: Illustration of the graph neural network architecture designed for the DR disease staging task. The architecture takes a heterogeneous graph as input and predicts the disease stage for the provided input graph. We generalize homogeneous message-passing architectures to the heterogeneous case by learning individual message-passing functions for each node type. After the message passing layers and linear layers, an aggregation for each node type is performed. Finally, the embeddings for all node types are aggregated and passed on to a classification head.
Our proposed network is restricted in its complexity compared to current CNNs and transformer models. Table 1 shows a comparison of the number of parameters and the training time of the GNN model and other prominent network architectures. Notably, the graph construction step is a prerequisite for our approach. We observe that with non-runtime-optimized code, the construction takes multiple seconds. Therefore, the benefit of faster training during hyperparameter tuning outweighs the construction overhead.
Table 1: Comparison of model complexities and training times for 100 epochs with a fixed batch size of 16 on a single NVIDIA RTX A6000. The heterogeneous GNN is more than 100 times smaller than the smallest evaluated CNN model. The * indicates that the model is fine-tuned without adapting all learnable weights. Pretraining of the RetFound model required around two weeks with eight NVIDIA Tesla A100 [zhou2023foundation].
@@
3.3 Interpretability Framework for Retinal Graphs
Our explainability framework consists of two components. First, an attribution method, and second, a localization method.
3.3.1 Attribution Method
We utilize integrated gradients [sundararajan2017axiomatic, mccloskey2019using] as attribution method to describe the impact of specific nodes and their characteristics on the algorithmic outcome, following the evaluations by Agarwal et al. [agarwal2023evaluating]. Alternatively, the explanations can be easily adjusted to any desired attribution method. The integrated gradients approach follows a straight line path in the input feature space from a baseline sample to the target sample. In computer vision, this baseline sample is typically a zero-image. Equivalent to baseline images, graph-based approaches require a baseline vector for each node in the graph. The attribution score $a_{v_{i}}$ of a feature $i$ of a node $v$ given a baseline $x_{bl}$ is then defined as